OpenStudy (anonymous):

What is the approximate area of the shaded region? A. 178.5 square centimeters B. 100 square centimeters C. 78.5 square centimeters D. 21.5 square centimeters

OpenStudy (anonymous):

Hello well is there a picture?

OpenStudy (anonymous):

Do you have a chart or a graph or something provided?????

OpenStudy (anonymous):

We can't answer it if you don't have anything to give your question to! :) @All(*)Dat

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Well, since you forgot to attach the picture I just googled the question with the answers and found this old post: http://openstudy.com/study#/updates/532ce1dde4b0593075e98ef5 In case that is indeed your question then you can tell that the circle is bounded by the square. Therefore we can tell that since the square's side is 10, then so is the circle's diameter. That means that the radius is half of 10, and therefore is 5. The area of circle is given by \(\pi r^2\), in our case \(\pi 5^2 = \pi 25 = 78.5398163397_{cm^2}\). That is approximately \(78.5_{cm^2}\)

OpenStudy (anonymous):

@All(*)Dat Send @Savanahdawson a message please. Apparently you blocked her for somehow. She just wants to know why. Thanks